A uniform electric field of 20 N/C exists along the x axis in space .the potential difference (VA-VB) for the points A(4m,2m) and B(6m,5m) is ____

Grade:12

2 Answers

Vikas TU
7 years ago
As the uniform Electric field lies only in along the x – axis the potential difference across two points A and B is:
V = Va – Vb = 20*2 = 40  V
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
As the uniform Electric field lies only in along the x – axis the potential difference across two points A and B is:
V = Va – Vb = 20*2 = 40  V
 
Thanks and regards,
Kushagra